### v3

@@ 3.1 Data Collection
MedQARo is obtained through the manual extraction of information from medical documents, namely medical summaries of patients that are registered either in the Oncology Department of Colţea Clinical Hospital, or in the “Dan Furtună” Medical Center, both located in Bucharest, Romania. The epicrises, originally drafted in Word format by the physicians attending each patient, were manually reviewed by a group of seven physicians involved in constructing MedQARo. In this group, two experienced physicians from Colţea Clinical Hospital, who led the data collection process, established the most prevalent cancer sites in their hospital, namely breast and lung. They selected patients with either one of these diagnoses, who received treatment at the hospital between 2019 and 2024. This selection process led to a set of 796 patients with breast cancer and 215 patients with lung cancer. To create a diverse multi-center and multi-cancer benchmark, 231 patients with five cancer types (head and neck, melanoma, renal cell carcinoma, bladder cancer, hepatocellular carcinoma) were included in the study. These patients are enrolled in cancer follow-up care programs at “Dan Furtună” Medical Center.
@@
The two experienced physicians further proposed a set of reference questions for each cancer site, comprising 48 questions for breast cancer, 61 questions for lung cancer, and 7 general questions that cover all other types. The other five physicians proposed 8 to 10 equivalent reformulations of the reference questions. Next, the physicians proceeded by extracting the answers for each unique question from each epicrisis and entering the respective answers into a database. To optimize annotation time, the questions for each epicrisis are grouped and annotated in the same batch. This means that a physician could read an epicrisis once, before answering multiple questions about the respective patient.
@@
Regarding research ethics, we emphasize that all data was fully anonymized prior to processing by removing any direct identifiers (names, dates of birth, personal identification numbers, etc.). The use of anonymized data for research purposes is based on the informed consent of patients, as recorded in the medical charts of Colţea Clinical Hospital and “Dan Furtună” Medical Center, and signed individually by each patient at admission. The database was approved by the Ethics Committee of Colţea Clinical Hospital where the research was conducted, in accordance with Decision No. 19091, dated October 5th, 2021. The database was also approved by the Ethics Committee of “Dan Furtună” Medical Center, as per Decision No. 1323, dated 11th December, 2025. The entire process complies with the requirements of the General Data Protection Regulation (GDPR), the principles of the Declaration of Helsinki, and the standards of good medical and research practices.
Table 2: A summary of the number of patients and corresponding QA pairs included in the MedQARo benchmark across cancer types. For breast and lung cancer, the splits are performed at the patient level, into training (70%), validation (15%), and in-domain test (15%), to prevent information leakage across splits. A cross-domain test set comprises a separate set of 231 patients with five cancer types (excluding breast and lung), from a different medical center. We report the number of patients per cancer type per split, and the corresponding number of QA samples generated for each split. The final row aggregates the total number of unique patients and QA pairs across different cancer types for each split. MedQARo comprises a total 105,880 QA pairs for 1,242 patients.
@@ 3.2 Preprocessing
Breast cancer. We collected data from 796 patients, each associated with a complete epicrisis, i.e. a detailed clinical summary for each patient, and a corresponding set of answers to 48 unique medical questions. To simulate natural variation in clinical queries, each question is reformulated by multiple physicians, leading multiple alternative formulations. For every patient-question pair, we randomly sample two question alternatives and generate two QA instances, each structured as a tuple of the form (patient ID, epicrisis, question, answer). This process resulted in a total of 76,416 unique examples.
Lung cancer. We compiled data from another 215 patients, each accompanied by an epicrisis summarizing their clinical history. For each patient, we curated a set of answers to 61 unique and clinically relevant questions. Similar to the breast cancer setup, each question is manually paraphrased into multiple linguistically diverse formulations to better reflect natural question variability. For each patient-question pair, we randomly sampled two different question variants. This leads to a total of 26,230 unique QA instances, capturing both clinical diversity and linguistic variation.
Other cancers. To construct a benchmark that enables the assessment of generalization capacity across medical centers and cancer types, we created a secondary dataset, which is not used during training. This dataset comprises 231 patients, diagnosed with a different kind of cancer (head and neck, melanoma, renal cell carcinoma, bladder cancer, or hepatocellular carcinoma), all coming from a different center than the patients included in our main dataset. For these patients, there are 7 general questions, relevant for all five cancer types. As for breast and lung cancer patients, we randomly sampled two different question variants for each patient-question pair, leading to a total of 3,234 unique QA pairs.
3.3 Partitioning and Statistics
@@
Figure 2: Distribution of question types in the MedQARo dataset. The dataset comprises three main categories: binary (yes/no) questions, extractive questions (answers explicitly found in the epicrisis), and reasoning questions (requiring inference beyond explicit mentions). Percentages are computed over the full set of 105,880 QA pairs. Best viewed in color.
@@
(a) Age distribution for breast cancer.
(b) Gender distribution for breast cancer.
(c) Age distribution for lung cancer.
(d) Gender distribution for lung cancer.
(e) Age distribution for other cancers.
(f) Gender distribution for other cancers.
Figure 3: Age and gender distributions of patients included in MedQARo. Rows correspond to breast cancer, lung cancer, and other cancers, respectively. The first column shows age distributions using 5-year bins (between 20 and 94 years), while the second column row shows gender distributions. Best viewed in color.
@@
3.4 Demographics
@@
Figure 4: Distribution across cancer types for patients included in the cross-domain test set of MedQARo. Best viewed in color.
@@ 4 Methods
4.1 Romanian-Adapted LLMs
@@
RoLLaMA2-7B-Instruct. As it inherits the architectural constraints of the original LLaMA 2 model [touvron2023llama], RoLLaMA2-7B is limited to a context window of 4,096 tokens. We employ a LoRA configuration specialized on the causal language modeling task. The chosen configuration enables updating a small subset of the model parameters ($0.0622\%$) during fine-tuning, significantly reducing training overhead. To enhance the adaptability of the LLM to the clinical QA task, we further consider several configuration alternatives. First, we evaluate on various prompt lengths, considering limits of 1,024, 2,048 and 4,096 tokens. If an epicrisis is longer than the accepted input length, we simply trim it from the end. When evaluating on prompts of 4,096 tokens, we also explore two prompt formats: Q+E+A (question + epicrisis + answer) and E+Q+A (epicrisis + question + answer). The model is trained via token-level cross-entropy, where the task is to predict the answer tokens via next token prediction. For this reason, the answer is always placed at the end of the prompt.
RoMistral-7B-Instruct. Given the large number of parameters in the base Mistral model, we again employ a LoRA-based fine-tuning strategy, enabling updates to only a small number of parameters. In this case, we fine-tune just $0.0470\%$ of the parameters, using the same LoRA configuration as for RoLLaMA2-7B-Instruct. This approach ensures both memory efficiency and effective domain adaptation to the Romanian clinical QA setting. We conduct experiments with two prompt lengths, namely 2,048 and 4,096. Consistent with the preliminary empirical results obtained with the previous model, the prompt format is set to Q+E+A in all the experiments performed with RoMistral-7B-Instruct. We employ token-level cross-entropy to optimize the model.
4.2 Long-Context LLM
@@ 4.4 Baselines
To assess the effectiveness of our fine-tuned models and to better isolate the impact of domain and task adaptation, we include four baselines.
@@
Random token selector. For this baseline, we run three versions with different input lengths (1,024, 2,048, and 4,096 tokens). For each setup, we randomly select the starting token index (between 0 and the maximum input length) and a span length (between 1 and 3 tokens) to serve as the predicted answer. The answer is extracted from the input context (epicrisis), while entirely ignoring the question. This baseline serves as a trivial reference point, highlighting the complexity of the clinical QA task and providing a lower-bound performance estimate.
@@
Majority answer model. Another simple baseline is based on answering with the most prevalent answer in the training set. This model consistently answers with “Nu” (“No”), regardless of the posed question or the epicrisis content. While still trivial, the majority answer model can be a more competitive baseline than the random token selector.
@@
Zero-shot inference. For each of the evaluated models (RoLLaMA2-7B-Instruct, RoMistral-7B-Instruct, Phi-4-mini-instruct, and LLaMA3-OpenBioLLM-8B), we assess performance without any task-specific fine-tuning. We use the Q+E+A prompt format across all models. The zero-shot setup is meant to reflect the out-of-the-box capabilities of distinct types of LLMs.

### v4

@@ Methods
MedQARo is obtained through the manual extraction of information from medical documents, namely medical summaries of patients that are registered either in the Oncology Department of Colţea Clinical Hospital, or in the “Dan Furtună” Medical Center, both located in Bucharest, Romania. The epicrises, originally drafted in Word format by the physicians attending each patient, were manually reviewed by a group of seven physicians involved in constructing MedQARo. In this group, two experienced physicians from Colţea Clinical Hospital, who led the data collection process, established the most prevalent cancer sites in their hospital, namely breast and lung. They selected patients with either one of these diagnoses, who received treatment at the hospital between 2019 and 2024. This selection process led to a set of 796 patients with breast cancer and 215 patients with lung cancer. To create a diverse multi-center and multi-cancer benchmark, 231 patients with one of five cancer types (head and neck, melanoma, renal cell carcinoma, bladder cancer, hepatocellular carcinoma) were included in the study. These patients are enrolled in cancer follow-up care programs at “Dan Furtună” Medical Center.
@@
The two experienced physicians further proposed a set of reference questions for each cancer site, comprising 48 questions for breast cancer, 61 questions for lung cancer, and 7 general questions that cover all other types. The other five physicians proposed 8 to 20 equivalent reformulations of the reference questions. Next, the physicians proceeded by extracting the answers for each unique question from each epicrisis and entering the respective answers into a database. To optimize annotation time, the questions for each epicrisis are grouped and annotated in the same batch. This means that a physician could read an epicrisis once, before answering multiple questions about the respective patient.
@@
Ethics Approval
Regarding research ethics, we emphasize that all data was fully anonymized prior to processing by removing any direct identifiers (names, dates of birth, personal identification numbers, etc.). The use of anonymized data for research purposes is based on the informed consent of patients, as recorded in the medical charts of Colţea Clinical Hospital and “Dan Furtună” Medical Center, and signed individually by each patient at admission. The database was approved by the Ethics Committee of Colţea Clinical Hospital (Anca Lupu, Ciprian Aldea, Traian Burcoş, Şerban Berteşteanu, Lavinia Coţofană, Letiţia Coriu, Gabriela Caragescu, Livia Neacşu) where the research was conducted, in accordance with Decision No. 19091, dated October 5th, 2021. The database was also approved by the Ethics Committee of “Dan Furtună” Medical Center (Bogdan Ioan Furtună, Cipriana Dimitrakopoulos, Alice Corina Fărcaş, Natalia Bărcaru, Daniela Tomescu), as per Decision No. 1323, dated 11th December, 2025. The entire process complies with the requirements of the General Data Protection Regulation (GDPR), the principles of the Declaration of Helsinki, and the standards of good medical and research practices.
Table 8: A summary of the number of patients and corresponding QA pairs included in the MedQARo benchmark across cancer types. For breast and lung cancer, the splits are performed at the patient level, into training (70%), validation (15%), and in-domain test (15%), to prevent information leakage across splits. A cross-domain test set comprises a separate set of 231 patients with one of five cancer types (excluding breast and lung), from a different medical center. We report the number of patients per cancer type per split, and the corresponding number of QA samples generated for each split. The final row aggregates the total number of unique patients and QA pairs across different cancer types for each split. MedQARo comprises a total 105,880 QA pairs for 1,242 patients.
@@
We collected data from 796 breast cancer patients, each associated with a complete epicrisis, i.e. a detailed clinical summary for each patient, and a corresponding set of answers to 48 unique medical questions. To simulate natural variation in clinical queries, each question is reformulated by multiple physicians, leading multiple alternative formulations. For every patient-question pair, we randomly sample two question alternatives and generate two QA instances, each structured as a tuple of the form (patient ID, epicrisis, question, answer). This process resulted in a total of 76,416 unique examples.
We compiled data from another 215 lung patients, each accompanied by an epicrisis summarizing their clinical history. For each patient, we curated a set of answers to 61 unique and clinically relevant questions. Similar to the breast cancer setup, each question is manually paraphrased into multiple linguistically diverse formulations to better reflect natural question variability. For each patient-question pair, we randomly sampled two different question variants. This leads to a total of 26,230 unique QA instances, capturing both clinical diversity and linguistic variation.
To construct a benchmark that enables the assessment of generalization capacity across medical centers and cancer types, we created a secondary dataset, which is not used during training. This dataset comprises 231 patients, diagnosed with a different kind of cancer (head and neck, melanoma, renal cell carcinoma, bladder cancer, or hepatocellular carcinoma), all coming from a different center than the patients included in our main dataset. For these patients, there are 7 general questions, relevant for all five cancer types. As for breast and lung cancer patients, we randomly sampled two different question variants for each patient-question pair, leading to a total of 3,234 unique QA pairs.
Data Partitioning and Statistics
@@
Figure 3: Distribution of question types in the MedQARo dataset. The dataset comprises three main categories: binary (yes/no) questions, extractive questions (answers explicitly found in the epicrisis), and reasoning questions (requiring inference beyond explicit mentions). Percentages are computed over the full set of 105,880 QA pairs. Best viewed in color. Blue represents extractive questions. Orange represents binary questions. Green represents reasoning questions.
Figure 4: Age and gender distributions of patients included in MedQARo. Separate distributions are reported for breast, lung, and other cancer patients. Age distributions are based on 5-year bins (between 20 and 94 years). a) Age distribution for breast cancer. b) Gender distribution for breast cancer. c) Age distribution for lung cancer. d) Gender distribution for lung cancer. e) Age distribution for other cancers. f) Gender distribution for other cancers.
Figure 5: Distribution across cancer types for patients included in the cross-domain test set of MedQARo. Blue represents head and neck cancers. Orange represents melanoma. Gray represents renal cell carcinoma. Yellow represents bladder cancer. Green represents hepatocellular carcinoma.
Data Demographics
@@
Romanian-Adapted LLMs
@@
As it inherits the architectural constraints of the original LLaMA 2 model [touvron2023llama], RoLLaMA2-7B is limited to a context window of 4,096 tokens. We employ a LoRA configuration specialized on the causal language modeling task. The chosen configuration enables updating a small subset of the model parameters ($0.0622\%$) during fine-tuning, significantly reducing training overhead. To enhance the adaptability of the LLM to the clinical QA task, we further consider several configuration alternatives. First, we evaluate on various prompt lengths, considering limits of 1,024, 2,048 and 4,096 tokens. If an epicrisis is longer than the accepted input length, we simply trim it from the end. When evaluating on prompts of 2,048 tokens, we also explore two prompt formats: Q+E+A (question + epicrisis + answer) and E+Q+A (epicrisis + question + answer). The model is trained via token-level cross-entropy, where the task is to predict the answer tokens via next token prediction. For this reason, the answer is always placed at the end of the prompt.
Given the large number of parameters in the base Mistral model, we again employ a LoRA-based fine-tuning strategy, enabling updates to only a small number of parameters. In this case, we fine-tune just $0.0470\%$ of the parameters, using the same LoRA configuration as for RoLLaMA2-7B-Instruct. This approach ensures both memory efficiency and effective domain adaptation to the Romanian clinical QA setting. We conduct experiments with two prompt lengths, namely 2,048 and 4,096. Consistent with the preliminary empirical results obtained with the previous model, the prompt format is set to Q+E+A in all the experiments performed with RoMistral-7B-Instruct. We employ token-level cross-entropy to optimize the model.
Long-Context LLM
@@
To assess the effectiveness of our fine-tuned models and to better isolate the impact of domain and task adaptation, we include four baselines:
@@
The random token selector is a baselines that randomly selects tokens for the input epicrisis. For this baseline, we run three versions with different input lengths (1,024, 2,048, and 4,096 tokens). For each setup, we randomly select the starting token index (between 0 and the maximum input length) and a span length (between 1 and 3 tokens) to serve as the predicted answer. The answer is extracted from the input context (epicrisis), while entirely ignoring the question. This baseline serves as a trivial reference point, highlighting the complexity of the clinical QA task and providing a lower-bound performance estimate.
@@
The majority answer model is another simple baseline is based on answering with the most prevalent answer in the training set. This model consistently answers with “Nu” (“No”), regardless of the posed question or the epicrisis content. While still trivial, the majority answer model can be a more competitive baseline than the random token selector.
@@
For each of the open-source LLMs (RoLLaMA2-7B-Instruct, RoMistral-7B-Instruct, Phi-4-mini-instruct, and LLaMA3-OpenBioLLM-8B), we assess performance without any task-specific fine-tuning. We use the Q+E+A prompt format across all models. The zero-shot inference setup is meant to reflect the out-of-the-box capabilities of distinct types of LLMs.